OpenStudy (anonymous):

What is the slope of the line which passes through (2, 6) and (0, 2)?

OpenStudy (lgbasallote):

\[\text{slope} = \frac{y_2 - y_1}{x_2 - x_1}\] \[\text{slope} = \frac{2 - 6}{0-2}\] \[\text{slope} = \frac{-4}{-2}\] simplify now